OpenStudy (anonymous):

Solve: 2(x + 4) − (x + 1) = 10

OpenStudy (anonymous):

Use the distributive property of multiplication again, then you will get rid of the parentheses. However, remember that -(a+b) = -a -b

OpenStudy (anonymous):

just show me how it is done

OpenStudy (anonymous):

show me with the equation

OpenStudy (anonymous):

2*x + 2*4 - x - 1 = 10

OpenStudy (anonymous):

and after what

OpenStudy (anonymous):

what you got now?

OpenStudy (anonymous):

2x-x-x=2

OpenStudy (anonymous):

no that's not right

OpenStudy (anonymous):

2x-x-1-2

OpenStudy (anonymous):

2x-x-1=2

OpenStudy (anonymous):

2x - (2*4) - x - 1

OpenStudy (anonymous):

so basically 2x - 8 - x - 1 = 10

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now simplify that

OpenStudy (anonymous):

(2x - x) - 8 - 1 = 10

OpenStudy (anonymous):

what's 2 times something - 1 time that same something?

OpenStudy (anonymous):

If you have two chuck norrisses and you subtract one chuck norris from that, how many chuck norrisses do you have?

jimthompson5910 (jim_thompson5910):

2(x + 4) - (x + 1) = 10 2x + 8 - x - 1 = 10 x + 7 = 10 x = 10 - 7 x = 3